# Diffuson-Dominated Thermal Transport Crossover from Ordered to Liquid-like Cu3BiS3:The Negligible Role of Ion Hopping

### v1

@@ 2.2 Diffuson-dominanted thermal transport
To accurately capture the finite-temperature lattice dynamics, the anharmonic phonon frequencies of crystalline Cu3BiS3 are calculated using self-consistent phonon theory with bubble correction (SCPB) 30, 51 (see Figures 2(a) and 2(b)). Interestingly, from Figures 2(a) and 2(b), we find that in both the and Pnma phases, low-frequency phonons exhibit softening with increasing temperature, in significant contrast to previously reported results for other crystalline compounds, such as BaZrO3 26, TlInTe2 52, and YbFe4Sb12 53. Note that in strongly anharmonic materials, such as CsPbBr351, KNbO3and NaNbO354, the negative energy shift from the bubble diagram 51 is significant and imposes a non-nigligible impact on both the theoretical phase transition temperature and the phonon linewidth. Specifically, considering only the energy shift from quartic anharmonicity leads to low-frequency phonon hardening in both the and Pnma phases of Cu3BiS3 (see Figure S4 in the Supporting Information). Therefore, incorporating the full explicit anharmonic vibrational effects, including both cubic and quartic anharmonicity, enables a more accurate description of phonon softening, which significantly influences the lattice thermal conductivity.
@@
With the finite-temperature phonons obtained, we proceed to calculate the lattice thermal conductivity of Cu3BiS3 using the Wigner transport equation 27, 28, accounting for both phonon population and coherence contributions, as shown in Figure 2(c). Our results show that the thermal conductivity of the phase at 400 K reaches 0.36 Wm-1K-1, in good agreement with the experimental value of 0.34 Wm-1K-1 17, thereby validating our calculations. More specifically, at 400 K, the coherence contribution to the total thermal conductivity is 0.22 Wm-1K-1, accounting for 6̃1% of the total thermal conductivity, highlighting the dominant role of phonon wave-like tunneling and strong lattice anharmonicity 55. We next continue to calculate the thermal conductivity of the ordered Pnma phase of Cu3BiS3 above 400 K using the Wigner transport equation 27, 28(see in Figure 2(c)). Our theoretical model predicts an ultra-low total thermal conductivity of 0.36 Wm-1K-1, comprising 0.16 Wm-1K-1 from the population contribution and 0.20 Wm-1K-1 from the coherence contribution. Surprisingly, the predicted thermal conductivity of the ordered Pnma phase of Cu3BiS3 agrees well with the experimentally observed value of 0.34 Wm-1K-1 at 400 K for the partially occupied Pnma phase. Furthermore, as temperature increases, the theoretical predictions remain in good agreement with experimental measurements for partial-occupied Cu3BiS3 (Pnma space group), as shown in Figure 2(c). The cross-phase thermal transport in Cu3BiS3 remains unexplored 16, motivating us to compare the predicted thermal conductivities of both the and Pnma phases with experimental values, as shown in the inset of Figure 2(c). Our results show that the thermal conductivity of the phase increases sharply above 400 K, deviating from the experimental values. In contrast, the Pnma phase shows a continuous decrease in thermal conductivity above 400 K, showing a good agreement with experimental results at high temperatures. These results further confirm the temperature-induced phase transition from to Pnma and provide a clear explanation for the smooth change in thermal conductivity at the transition point.
@@
Considering the liquid-like behavior of the Pnma phase at high temperatures, we calculate using the Green-Kubo equilibrium molecular dynamics (GK-EMD) method 38 (see Figure S5 in the Supporting Information), which accounts for all terms in the heat flux operator 39, 40. As shown in Figure 2(c), the total thermal conductivity predicted by the GK-EMD method at 400 K is in excellent agreement with that obtained from the Wigner transport equation 27, 28. Even at 500 K, the discrepancy between the two methods remains negligible, with a difference of only 0.01 Wm-1K-1. Based on the above observation, and considering that the thermal conductivity from the Wigner transport equation 27, 28 is calculated for the ordered Pnma phase—i.e., in the absence of ion hopping—we conclude that ion hopping has minimal impact on thermal transport in liquid-like Cu3BiS3. The negligible effect of ion hopping on thermal transport can be attributed to the fact that ion-hopping behavior resembles rattling-like motion observed in certain crystalline materials 13, 56, which induces strong anharmonicity. This, in turn, suppresses phonon transport while enhancing the coherence contribution. Thus, in diffuson-dominated thermal transport regimes, ion hopping contributes minimally to the total thermal conductivity.
@@
To further validate our theoretically predicted lattice thermal conductivity of crystalline Cu3BiS3, we compared the calculated values with experimental measurements in the temperature range of 2 to 100 K (see Figure 2(d)). Clearly, when boundary scattering at low temperatures is not considered 18, the predicted thermal conductivity—accounting only for three-phonon, four-phonon, and phonon-isotope scattering—deviates significantly from the experimental values. To accurately reproduce the experimental thermal conductivity 57, 58, effective grain sizes of 50, 100, and 200 nm are introduced, which significantly suppress the thermal conductivity and bring the predictions closer to the experimental results. On top of the 200 nm boundary scattering model, the inclusion of additional point-defect scattering leads to improved agreement between theory and experiment, as depicted in Figure 2(d). In addition, we performed a comparative analysis between the experimentally measured isobaric heat capacity (Cp) of Cu3BiS3 and the corresponding theoretical predictions within the temperature range of 2–70 K (see Figure S6 in the Supporting Information). Notably, at low temperatures, the experimental Cp values exhibit excellent consistency with the calculated constant volume heat capacity (Cv), reflecting the reliability of the theoretical model in this regime. Meanwhile, our computational results closely reproduce the isobaric heat capacity data reported by Wu et al. prior to the observed phase transition 16.
2.3 Microscopic mechanisms of thermal transport
@@
Next, to reveal the microscopic mechanisms of thermal transport in crystalline Cu3BiS3 with the and phases, we calculate the spectral and cumulative propagating thermal conductivity at 300 K and 400 K, respectively. As shown in Figure 3(a), phonons with frequencies below 1.5 THz dominate the particle-like phonon channel, which can be attributed to the relatively large group velocities of the acoustic and low-frequency optical modes (see Figure S7 in the Supporting Information). To elucidate the correlation between the ultra-low propagating thermal conductivity and the atomic structure of crystalline Cu3BiS3, we project the atomic contributions onto the phonon bands, as shown in Figure 3(b). Specifically, Cu atoms predominantly contribute to the low- and mid-frequency optical phonon modes in the range of 1.5–5.0 THz. In contrast, S atoms primarily participate in the high-frequency optical modes above 5.0 THz, while Bi atoms exhibit relatively low participation and are mainly involved in a limited number of low-frequency optical modes between 1.5 and 2.0 THz (see Figure S8 in the Supporting Information). The fully Cu- and Bi-occupied low-frequency optical modes lead to strong phonon scattering rates, thereby suppressing thermal transport in crystalline Cu3BiS3, a phenomenon similarly observed in CsCu4Se3 55. Importantly, Cu atoms give rise to low-energy, Einstein-like quasilocalized vibrational modes 56, particularly within the 2–4 THz frequency range. These Cu-dominated quasilocalized vibrational phonon modes act as strong phonon scatterers, significantly interacting with and disrupting the propagation of heat-carrying acoustic phonons. This phenomenon has also been reported in cubic SrTiO359, the double perovskite Cs2AgBiBr630, and copper-based chalcogenides such as o-CsCu5S324.
@@
As illustrated in Figure 3(c), we further calculate the phonon lifetimes for the phase at 300 K and the phase at 400 K, respectively. We find that the majority of phonon modes exhibit lifetimes shorter than the Wigner limit 27, 28, considering both three-phonon (3ph) and four-phonon (4ph) scattering processes, indicating strong anharmonicity in the system. For the phase, we observe a dip in phonon lifetimes for phonons around 0.75 THz, which are dominated by Bi atoms, highlighting a strong scattering source that contributes to the suppression of thermal conductivity 26, 30. Similarly, the flattened, localized vibrational mode, primarily associated with Cu atoms near 2.2 THz, also exhibits an extremely short lifetime and pronounced anharmonic scattering (see Figure 3(c) and Figure S9 in the Supporting Information). Therefore, we further confirm that the ultra-low thermal conductivity is attributed to the vibrations of Cu and Bi atoms. In the phase, Cu atoms play a similar role, as the Cu-dominated mode near 1.8 THz exhibits a significant drop in phonon lifetime (see Figure 3(c) and Figure S9 in the Supporting Information), corresponding to the dip observed in the spectral thermal conductivity (see Figure 3(a)). Although strong anharmonicity and the low phonon lifetimes of quasilocalized modes suppress particle-like phonon propagation 25, they enhance the wave-like phonon diffuson contribution (see Figure 3(d)). Theoretically, the presence of dense, quasilocalized phonon modes with similar energies and strong anharmonicity can facilitate significant heat exchange between phonon modes in energy space 60, 55, thereby promoting diffusion-like transport 61.
@@
To gain a better understanding of microscopic mechanisms behind the coherences’ thermal conductivity in Cu3BiS3, we calculate the spectral and cumulative mode-specific contributions to coherences’ conductivity at 300 and 400 K, respectively, as shown in Figure 3(d). In sharp contrast to the spectral populations’ conductivity of Cu3BiS3 in Figure 3(b), the majority of the coherences’ conductivity from the wavelike tunneling channel is carried by the phonons with almost full-frequency at both 300 and 400 K. This can be attributed to the small inter-band spacing within full-frequency region (due to dense phonon dispersions (see Figure S10 in the Supporting Information)) in conjunction with large linewidths (large scattering rates (see Figure 3(c)) or strong anharmonicity. Specifically, the spectral coherence conductivity exhibits a pronounced peak at 2.2 THz in the phase and at 1.8 THz in the Pnma phase. These peaks correspond to Einstein-like quasilocalized vibrational modes predominantly associated with Cu atoms (see Figure 3(b)). Within the middle and high-frequency domains, the contribution from spectral coherences’ conductivity is considerable, largely driven by strong anharmonic scattering alongside small interband spacing. Overall, for the Cu3BiS3,strong anharmonic effects combined with reduced phonon group velocities lead to quasilocalization in real space, thereby suppressing phonon-gas transport and fostering strong phonon coupling alongside significant diffuson contributions. Figures 4(a-f) show that the contributions to coherences’ thermal conductivity in Cu3BiS3 at 300 and 400 K, can be resolved in terms of the phonon energies, namely, and of two coupled phonons. At 300 K, similarly to the harmonic glasses 62, where quasi-degenerate phonon states dominate (see Figures 4(a-c)), the phonons contributing to along the three axis have relatively the same frequencies. For the pnma phase at 400 K, the two-dimensional density of states for the coherences’ thermal conductivity along the -aixs and -aixs (see Figures 4(d-e)) also shows couplings between quasi-degenerate states. In contrast, the the coherences’ thermal conductivity along the -aixs involves couplings between phonon modes with significantly different frequencies (see Figure 4(f)), driven by strong anharmonicity. As a result, the associated heat transport mechanism is fundamentally different from that in harmonic glasses.
@@ 3 Conclusion
In summary, we have comprehensively investigated the anharmonic lattice dynamics and thermal transport behavior in both ordered and partially disordered Cu3BiS3 using a combination of experimental measurements and advanced DFT-based theoretical simulations. To begin with, we experimentally synthesized and characterized both the ordered and liquid-like, partially Cu-occupied structures of Cu3BiS3, with the Pnma phase emerging upon temperature increase-an observation further supported by ab initio molecular dynamics simulations. Subsequently, we employ DFT-based self-consistent phonon theory with bubble-diagram corrections, combined with the Wigner transport equation, to capture the anharmonic lattice dynamics and uncover the microscopic mechanisms underlying the ultra-low thermal conductivity in Cu3BiS3. At low temperatures (2-100 K), the predicted thermal conductivity-accounting for 3ph, 4ph, isotope, 200 nm boundary, and point-defect scattering, accurately reproduce the experimental thermal conductivity. At 400 K, our results demonstrate that both the ordered phase and the partially Cu-occupied phase exhibit ultralow thermal conductivity, with values of 0.36 Wm-1K-1. Notably, our predicted thermal conductivity, incorporating both propagating and diffuson-like phonon contributions, well reproduces the experimental measurements of 0.34 Wm-1K-1, even for the partially Cu-occupied phase. In both the and phases, the ultralow thermal conductivity is mainly due to strong anharmonic scattering associated with Bi and Cu atoms. In particular, strong antibonding states between Cu and Se bonds give rise to the weak bonding of these Cu atoms, which leads to low-energy, Einstein-like vibrational modes that efficiently scatter heat-carrying acoustic phonons, thereby hindering their particle-like propagation. On the other hand, machine-learning-based Green–Kubo molecular dynamics simulations further confirm that ion hopping has a negligible impact on heat transport, even in the high-temperature, liquid-like phase. Therefore, despite the disordered nature of Cu3BiS3, our models—both machine learning and first-principles-based—provide reasonable estimates of thermal conductivity when using ordered structural representations. Our findings on Cu3BiS3 suggest that the lattice thermal conductivity () of amorphous or disordered compounds can, to a reasonable degree of accuracy, be approximated using calculations based on their ordered crystalline counterparts. This approach provides a practical pathway for evaluating thermal transport properties in complex systems where fully capturing structural disorder remains computationally challenging. Moreover, our study highlights the minimal impact of ionic mobility on thermal conductivity and offers valuable insights into the fundamental phonon transport mechanisms underlying the ultralow observed in both the ordered and liquid-like phases of Cu3BiS3.

### v2

@@ 3 Diffuson-dominated thermal transport
To accurately capture the finite-temperature lattice dynamics, the anharmonic phonon frequencies of crystalline Cu3BiS3 are calculated using self-consistent phonon theory with bubble correction (SCPB) [32, 53] (see Figures 3(a) and 3(b)). Interestingly, from Figures 3(a) and 3(b), we find that in both the and phases, low-frequency phonons exhibit softening with increasing temperature, in significant contrast to previously reported results for other crystalline compounds, such as BaZrO3 [27], TlInTe2 [54], and YbFe4Sb12 [55]. Note that in strongly anharmonic materials, such as CsPbBr3 [53], KNbO3 and NaNbO3 [56], the negative energy shift from the bubble diagram [53] is significant and imposes a non-negligible impact on both the theoretical phase transition temperature and the phonon linewidth. Specifically, considering only the energy shift from quartic anharmonicity leads to low-frequency phonon hardening in both the and phases of Cu3BiS3 (see Figure S5 in the Supporting Information). Therefore, incorporating the full explicit anharmonic vibrational effects, including both cubic and quartic anharmonicity, enables a more accurate description of phonon softening, which significantly influences the lattice thermal conductivity.
@@
With the finite-temperature phonons obtained, we proceed to calculate the lattice thermal conductivity of Cu3BiS3 using the Wigner transport equation [28, 29], accounting for both phonon population and coherence contributions, as shown in Figure 3(c). Our results show that the thermal conductivity of the phase at 400 K reaches 0.36 Wm-1K-1, in good agreement with the experimental value of 0.34 Wm-1K-1 [17], thereby validating our calculations. More specifically, at 400 K, the coherence contribution to the total thermal conductivity is 0.22 Wm-1K-1, accounting for 6̃1% of the total thermal conductivity, highlighting the dominant role of phonon wave-like tunneling and strong lattice anharmonicity [57]. We next continue to calculate the thermal conductivity of the phase of Cu3BiS3 above 400 K using the Wigner transport equation [28, 29](see in Figure 3(c)). Our theoretical model predicts an ultra-low total thermal conductivity of 0.36 Wm-1K-1, comprising 0.16 Wm-1K-1 from the population contribution and 0.20 Wm-1K-1 from the coherence contribution. Surprisingly, the predicted thermal conductivity of the phase of Cu3BiS3 agrees well with the experimentally observed value of 0.34 Wm-1K-1 at 400 K for the partially occupied phase. Furthermore, as temperature increases, the theoretical predictions remain in good agreement with experimental measurements for partial-occupied Cu3BiS3 ( space group), as shown in Figure 3(c). The cross-phase thermal transport in Cu3BiS3 remains unexplored [16], motivating us to compare the predicted thermal conductivities of both the and phases with experimental values, as shown in the inset of Figure 3(c). Our results show that the thermal conductivity of the phase increases sharply above 400 K, deviating from the experimental values. In contrast, the phase shows a continuous decrease in thermal conductivity above 400 K, showing a good agreement with experimental results at high temperatures. These results further confirm the temperature-induced phase transition from to and provide a clear explanation for the smooth change in thermal conductivity at the transition point. It is worth noting that the Wigner transport equation is limited by its reliance on the anharmonic heat flux contributions [29], and higher-order phonon scattering. Moreover, it cannot capture liquid-like ion diffuson [58], such as the motion of ions in disordered or partially liquid environments. In contrast, molecular dynamics methods naturally include these effects, making them more suitable for such scenarios.
@@
Considering the liquid-like behavior of the phase at high temperatures, we calculate using the Green-Kubo equilibrium molecular dynamics (GK-EMD) method [40] (see Figure S6 in the Supporting Information), which accounts for all terms in the heat flux operator [41, 42]. As shown in Figure 3(c), the total thermal conductivity predicted by the GK-EMD method at 400 K is in excellent agreement with that obtained from the Wigner transport equation [28, 29]. Even at 500 K, the discrepancy between the two methods remains negligible, with a difference of only 0.01 Wm-1K-1. Based on the above observation, and considering that the thermal conductivity from the Wigner transport equation [28, 29] is calculated for the phase—i.e., in the absence of ion hopping—we conclude that ion hopping has minimal impact on thermal transport in liquid-like Cu3BiS3. The negligible effect of ion hopping on thermal transport can be attributed to the fact that ion-hopping behavior resembles rattling-like motion observed in certain crystalline materials [13, 59], which induces strong anharmonicity. This, in turn, suppresses phonon transport while enhancing the coherence contribution. Thus, in diffuson-dominated thermal transport regimes, ion hopping contributes minimally to the total thermal conductivity.
@@
To further validate our theoretically predicted lattice thermal conductivity of crystalline Cu3BiS3, we compared the calculated values with experimental measurements in the temperature range of 2 to 100 K (see Figure 3(d)). Clearly, when boundary scattering at low temperatures is not considered [18], the predicted thermal conductivity—accounting only for three-phonon, four-phonon, and phonon-isotope scattering—deviates significantly from the experimental values. To accurately reproduce the experimental thermal conductivity [60, 61], effective grain sizes of 50, 100, and 200 nm are introduced, which significantly suppress the thermal conductivity and bring the predictions closer to the experimental results. On top of the 200 nm boundary scattering model, the inclusion of additional point-defect scattering leads to improved agreement between theory and experiment, as depicted in Figure 3(d). In addition, we performed a comparative analysis between the experimentally measured isobaric heat capacity (Cp) of Cu3BiS3 and the corresponding theoretical predictions within the temperature range of 2–70 K (see Figure S7 in the Supporting Information). Notably, at low temperatures, the experimental Cp values exhibit excellent consistency with the calculated constant volume heat capacity (Cv), reflecting the reliability of the theoretical model in this regime. Meanwhile, our computational results closely reproduce the isobaric heat capacity data reported by Wu et al. prior to the observed phase transition [16].
4 Microscopic mechanisms of thermal transport
@@
Next, to reveal the microscopic mechanisms of thermal transport in crystalline Cu3BiS3 with the and phases, we calculate the spectral and cumulative propagating thermal conductivity at 300 K and 400 K, respectively. As shown in Figure 4(a), phonons with frequencies below 1.5 THz dominate the particle-like phonon channel, which can be attributed to the relatively large group velocities of the acoustic and low-frequency optical modes (see Figure S8 in the Supporting Information). To elucidate the correlation between the ultra-low propagating thermal conductivity and the atomic structure of crystalline Cu3BiS3, we project the atomic contributions onto the phonon bands, as shown in Figure 4(b). Specifically, Cu atoms predominantly contribute to the low- and mid-frequency optical phonon modes in the range of 1.5–5.0 THz. In contrast, S atoms primarily participate in the high-frequency optical modes above 5.0 THz, while Bi atoms exhibit relatively low participation and are mainly involved in a limited number of low-frequency optical modes between 1.5 and 2.0 THz (see Figure S9 in the Supporting Information). The fully Cu- and Bi-occupied low-frequency optical modes lead to strong phonon scattering rates, thereby suppressing thermal transport in crystalline Cu3BiS3, a phenomenon similarly observed in CsCu4Se3 [57]. Importantly, Cu atoms give rise to low-energy, Einstein-like quasilocalized vibrational modes [59], particularly within the 2–4 THz frequency range. These Cu-dominated quasilocalized vibrational phonon modes act as strong phonon scatterers, significantly interacting with and disrupting the propagation of heat-carrying acoustic phonons. This phenomenon has also been reported in cubic SrTiO3[62], the double perovskite Cs2AgBiBr6[32], and copper-based chalcogenides such as o-CsCu5S3[24].
@@
As illustrated in Figure 4(c), we further calculate the phonon lifetimes for the phase at 300 K and the phase at 400 K, respectively. We find that the majority of phonon modes exhibit lifetimes shorter than the Wigner limit [28, 29], considering both three-phonon (3ph) and four-phonon (4ph) scattering processes, indicating strong anharmonicity in the system. For the phase, we observe a dip in phonon lifetimes for phonons around 0.75 THz, which are dominated by Bi atoms, highlighting a strong scattering source that contributes to the suppression of thermal conductivity [27, 32]. Similarly, the flattened, localized vibrational mode, primarily associated with Cu atoms near 2.2 THz, also exhibits an extremely short lifetime and pronounced anharmonic scattering (see Figure 4(c) and Figure S10 in the Supporting Information). Therefore, we further confirm that the ultra-low thermal conductivity is attributed to the vibrations of Cu and Bi atoms. In the phase, Cu atoms play a similar role, as the Cu-dominated mode near 1.8 THz exhibits a significant drop in phonon lifetime (see Figure 4(c) and Figure S10 in the Supporting Information), corresponding to the dip observed in the spectral thermal conductivity (see Figure 4(a)). Although strong anharmonicity and the low phonon lifetimes of quasilocalized modes suppress particle-like phonon propagation [25], they enhance the wave-like phonon diffuson contribution (see Figure 4(d)). Theoretically, the presence of dense, quasilocalized phonon modes with similar energies and strong anharmonicity can facilitate significant heat exchange between phonon modes in energy space [63, 57], thereby promoting diffuson-like transport [64].
@@
To gain a better understanding of microscopic mechanisms behind the coherences’ thermal conductivity in Cu3BiS3, we calculate the spectral and cumulative mode-specific contributions to coherences’ conductivity at 300 and 400 K, respectively, as shown in Figure 4(d). In sharp contrast to the spectral populations’ conductivity of Cu3BiS3 in Figure 4(b), the majority of the coherences’ conductivity from the wavelike tunneling channel is carried by the phonons with almost full-frequency at both 300 and 400 K. This can be attributed to the small inter-band spacing within full-frequency region (due to dense phonon dispersions (see Figure S11 in the Supporting Information)) in conjunction with large linewidths (large scattering rates (see Figure 4(c)) or strong anharmonicity. Specifically, the spectral coherence conductivity exhibits a pronounced peak at 2.2 THz in the phase and at 1.8 THz in the phase. These peaks correspond to Einstein-like quasilocalized vibrational modes predominantly associated with Cu atoms (see Figure 4(b)). Within the middle and high-frequency domains, the contribution from spectral coherences’ conductivity is considerable, largely driven by strong anharmonic scattering alongside small interband spacing. Overall, for the Cu3BiS3,strong anharmonic effects combined with reduced phonon group velocities lead to quasilocalization in real space, thereby suppressing phonon-gas transport and fostering strong phonon coupling alongside significant diffuson contributions. Figures 5(a-f) show that the contributions to coherences’ thermal conductivity in Cu3BiS3 at 300 and 400 K, can be resolved in terms of the phonon energies, namely, and of two coupled phonons. At 300 K, similarly to the harmonic glasses [65], where quasi-degenerate phonon states dominate (see Figures 5(a-c)), the phonons contributing to along the three axe have relatively the same frequencies. For the phase at 400 K, the two-dimensional density of states for the coherences’ thermal conductivity along the -aixs and -aixs (see Figures 5(d-e)) also shows couplings between quasi-degenerate states. In contrast, the the coherences’ thermal conductivity along the -aixs involves couplings between phonon modes with significantly different frequencies (see Figure 5(f)), driven by strong anharmonicity. As a result, the associated heat transport mechanism is fundamentally different from that in harmonic glasses.
@@ 5 Conclusion
In summary, we have comprehensively investigated the anharmonic lattice dynamics and thermal transport behavior in both ordered and partially disordered Cu3BiS3 using a combination of experimental measurements and advanced DFT-based theoretical simulations. To begin with, we experimentally synthesized and characterized both the ordered structure stable at room temperature and liquid-like structure occurring above 400 K of Cu3BiS3. The latter disordered structure crystallizing in the space group, is characterized by statistical distribution of Cu+ ions on several crystallographic sites and emerges upon temperature increase - an observation further supported by ab initio molecular dynamics simulations. Subsequently, we employ DFT-based self-consistent phonon theory with bubble-diagram corrections, combined with the Wigner transport equation, to capture the anharmonic lattice dynamics and uncover the microscopic mechanisms underlying the ultra-low thermal conductivity in Cu3BiS3. At low temperatures (2-100 K), the predicted thermal conductivity-accounting for 3ph, 4ph, isotope, 200 nm boundary, and point-defect scattering, accurately reproduce the experimental thermal conductivity. At 400 K, our results demonstrate that both the ordered phase and disordered phase exhibit ultralow thermal conductivity, with values of 0.36 Wm-1K-1. Notably, our predicted thermal conductivity, incorporating both propagating and diffuson-like phonon contributions, well reproduces the experimental measurements of 0.34 Wm-1K-1, even for the partially Cu-occupied phase. In both the and phases, the ultralow thermal conductivity is mainly due to strong anharmonic scattering associated with Bi and Cu atoms. In particular, large anisotropic vibrations of Cu in triangular coordination leads to low-energy, Einstein-like vibrational modes that efficiently scatter heat-carrying acoustic phonons, thereby hindering their particle-like propagation. On the other hand, machine-learning-based Green–Kubo molecular dynamics simulations further confirm that ion hopping has a negligible impact on heat transport, even in the high-temperature, liquid-like phase. Therefore, despite the disordered nature of Cu3BiS3, our models—both machine learning and first-principles-based—provide reasonable estimates of thermal conductivity when using ordered structural representations. Our findings on Cu3BiS3 suggest that the lattice thermal conductivity () of amorphous or disordered compounds can, to a reasonable degree of accuracy, be approximated using calculations based on their ordered crystalline counterparts. This approach provides a practical pathway for evaluating thermal transport properties in complex systems where fully capturing structural disorder remains computationally challenging. Moreover, our study highlights the minimal impact of ionic mobility on thermal conductivity and offers valuable insights into the fundamental phonon transport mechanisms underlying the ultralow observed in both the ordered and liquid-like phases of Cu3BiS3.
Acknowledgements